Find the additive inverse of the following value mod m.

16. 7 mod 9

17. 4 mod 12

18. 63 mod 30

19. 222 mod 3
Answer #1

Additive inverse means "additive inverse of a number a is the number that, when added to a, yields zero".

property a+(-a)=0 (where 'a' is a whole number)

16.

the value of 7 mod 9=7

so, the additive inverse of 7 mod 9 is -7 [7+(-7)=0]

17.

the value of 4mod12 is 4

so,the additive inverse of 4mod12 is -4 [4+(-4)=0]

18.

the value of 63 mod 30 is 3

so, the additive inverse of 63 mod 30 is -3 [3+(-3)=0]

19.

the value of 222 mod 3 is 0

so,the additive inverse of 222 mod 3 is 0